A stone is thrown from the top of a building TT toward another building 50m away. The initial velocity of the stone is 20m/s, at 400 above horizontal. How far above or below its original level will the ball strike the opposite wall?


1
Expert's answer
2021-01-31T03:41:59-0500

"v_{x0}=20\cdot\cos40°=15.32(m/s)"


"v_{y0}=20\cdot\sin40°=12.86(m/s)"


The stone flight time


"t=50\/15.32=3.264(s)"


"y=v_{y0}t-gt^2/2=12.86\cdot 3.264-9.8\cdot3.264^2/2=-10.23(m)"


"10.23\ (m)" below its original level. Answer